### v1

@@ 3.2.1 Existing Benchmarks:
Both supervised:MPD and supervised:MPD IB models are trained on ImageNet dataset and evaluated on existing benchmarks and ImageNet-PD. ImageNet-X [18] and ImageNet-E [22]) considered as they are partially affected by perspective distortion. ImageNet-E[22] is a recent benchmark dataset specifically designed to evaluate the robustness of image classifiers concerning object attributes, including background settings, object sizes, spatial positioning, and orientation. ImageNet-X[18] introduces a comprehensive set of sixteen human annotations focusing on attributes like pose, background, and lighting derived from the ImageNet.
Figure 6: ImageNet-E: Drop of Top1 accuracy under background changes, size changes, random position (rp), random direction (rd), and average over 11 subsets. Lower is better. The mean is reported for size and background-related subsets in ImageNet-E[22]. Average reports the mean of all subsets.
Figure 7: Activation maps of ImageNet-E examples. ’redbone’ from rd subset and ’peke’ from random direction (rd) subset.
@@ 3.2.2 Benchmarking on ImageNet-PD:
Table 1: Comparative analysis of MPD trained on supervised (supervised:MPD model) and self-supervised (ssl:MPD model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background, without background interpolation). The probability $P$ of applying MPD in model fine-tuning is in the first column. More results on probabilities 0.0 and 1.0 in Table 8 in suppl. material.
@@
Both supervised:MPD and supervised:MPD IB demonstrate notable improvements (Fig. 9 (a)-(b)) on ImageNet-PD subsets, achieving average Top-1 accuracy enhancements of 9.42% and 9.54% respectively, over the standard ResNet50 model. Please refer to Table 1 and Table 9 for detailed comparison; Table 9 is in suppl. material.
@@
Concurrently, these models maintain performance on the original ImageNet validation set, highlighting their efficacy and adaptability amidst perspectiveness. The activation maps shown in Fig. 5 (row 2) further corroborate this outperforming trend. Ablation on probability of applying MPD and MPD IB as a data augmentation (Fig. 9 (c)-(d)) exhibit consistent performance.
@@ 3.3 MPD’s effects on self-supervised learning
MPD is incorporated in SimCLR and DINO as additional data augmentation, as shown in Fig. 17 and 18 respectively. Both ssl:MPD and ssl:MPD IB models are trained on ImageNet dataset and evaluated on existing benchmarks and ImageNet-PD.
@@ 3.3.1 Existing Benchmarks:
ssl:MPD and ssl:MPD IB models demonstrate an enhancement in handling perspective-related variations in ImageNet-E as shown in Fig. 6. Specifically, it shows a reduction in the accuracy drop for position (1.21%), direction (1.41%), and size (0.48%) changes compared to the standard ResNet50, indicating its robustness against PD. On the ImageNet-X benchmark, ssl:MPD and ssl:MPD IB models outperform with an average accuracy of 62.99%, compared to the standard ResNet50’s 60.64%, and also show a reduced error ratio of 1.41 against the ResNet50’s 1.55 (refer Fig. 8). Notably, performance improvement is evident in managing variations in size and shape, showcasing its robustness against PD. Detailed results are in Table 11 in suppl. material.
@@ 3.3.2 Benchmarking on ImageNet-PD:
ssl:MPD model (Fig. 10 (a)) shows an average improvement of approximately 10.32% across the PD subsets, while the ssl:MPD IB (Fig. 10 (b)) demonstrates an average increase of around 10.14%. In addition, both variants display robust performance on the original ImageNet validation set. Please refer to Table 1 and Table 9 for detailed comparison; Table 9 is in suppl. material. The activation maps depicted in Fig. 5 (row 3) further validate the improved performance. Moreover, ablation on the probability of applying MPD and MPD IB as data augmentation (Fig. 10 (c)-(d)) demonstrates a consistent increase in performance. We have also performed linear evaluations on self-supervised method SimCLR [7] and DINO [6] (refer Table 14, 15 for detailed results; Fig. 20 for activation maps in suppl. material). Similarly, significant improvement is noticed on ImageNet-PD subsets while maintaining the performance on standard ImageNet validation set as shown in Fig. 11 (also refer Table 12, 13 for detailed results in suppl. material).
@@ 3.4 MPD’s generalizability
This section discusses MPD’s adaptability in different applications, namely crowd counting (CC), fisheye, and person re-identification, where perspectiveness is evident.
@@ 3.4.1 Crowd counting:
Numerous human heads within a single image present challenges due to the variations in size, orientation, and shape, which naturally includes perspectiveness. We investigate the potential of MPD in crowd counting on several publicly available datasets, including ShanghaiTech (SHTech) Part A and Part B[47] and UCF-CC50[16]. The ShanghaiTech Part A features highly congested scenes, and Part B consists of sparser scenes captured on a busy street. The UCF-CC50 dataset comprises 50 diverse scenes acquired from the internet, featuring a considerable variation in the crowd.
MPD-CC: MPD is adapted to CC by extending its capability to transform scene images along with their CC labels, as mentioned in steps 4-7 of algorithm 2.
MPD-AutoCrowd: We introduce MPD-AutoCrowd (Algorithm 2), a novel approach for the automated augmentation of crowd in scene images by upgrading MPD-CC. It seamlessly integrates synthetic crowds into the background regions of transformed images, complete with corresponding labels. This approach facilitates scene-level crowd augmentation without human supervision (Fig. 12).
@@
MPD-CC and MPD-AutoCrowd are integrated with P2P-Net[36] to evaluate the effectiveness of MPD in crowd counting. Training parameters, CNN architecture (VGG16), and evaluation configuration are used as mentioned in P2P-Net[36]. The probability of applying MPD-CC and MPD AutoCrowd is 0.1, and real and imaginary components of $c$ range from $0.2$ to $0.6$. This method has yielded improved results, surpassing previous state-of-the-art performance. Please refer to Table 2 for a detailed comparison with other CC methods. MPD-CC and MPD-AutoCrowd outperform on multiple benchmarks.
MPD-AutoCrowd has a MAE of 50.81 on SHTech Part A and 6.73 on SHTech Part B. The performance on the UCF-CC50 ( a highly dense crowds) dataset, where MPD-AutoCrowd has an MAE of 96.80 and a Mean Squared Error (MSE) of 139.50, significantly outperforms other models. We further improved results on UCF-CC50 with ResNet50 encoder as shown in Table 16 in suppl. material.
@@ 3.4.2 Transfer learning on fisheye images
Fisheye image distortions, caused by a wide field of view, are tackled using spherical perspective models as detailed in [48, 21]. MPD simulates spherical views in pixel space and improves representation learning for fisheye images. We showcase the transferable capability of models trained with MPD and MPD IB through supervised and self-supervised methods on the ImageNet dataset, applying them to the VOC-360 fisheye dataset[11]. This dataset encompasses 39,575 fisheye images with multi-label classification annotations. We fine-tune these models on the VOC360 dataset, running for 100 epochs with a batch size 16. We show the performance of supervised and self-supervised models for the multi-label classification task in Table 3. ResNet50 is used as backbone encoder. Ablations on label efficiency (Table 17, 18, and Fig. 21) are given in the suppl. material.
@@ 3.4.3 Person Re-Identification:
The DeepSportRadar re-identification dataset[38] includes video frames with varying poses and camera angles, offering training (436 queries, 8133 gallery images) and testing sets (50 queries, 910 gallery images) to address perspective variability in re-identification tasks. We incorporated the MPD within the state-of-the-art CLIP-ReIdent method[12]. The probability of applying MPD is 0.1, and real & imaginary components of $c$ range from $0.2$ to $0.6$. CLIP-ReIdent is CLIP[33]-adapted contrastive image-image pre-training for person re-identification, leveraging transformer architecture. MPD-incorporated Clip-ReIdent models improve performance with both backbone architectures, ViT-L-14 and ResNet50x16, as shown in Table 5. Ablations on MPD parameters are given in suppl. material in Table 19. In addition, it boosts performance across different frame counts, achieving competitive results even with limited frames and performance further enhanced by longer training, as shown in Table 5.
@@ 4 Conclusion
In this paper, we address perspective distortion for robust computer vision applications. The proposed MPD method models real-world distortions by applying a family of Möbius transforms on existing data without estimating camera parameters or relying on actual distorted data. We develop a perspectively-distorted ImageNet-PD dataset to benchmark the robustness of the deep learning models and reveal that existing models lack robustness. Further, MPD-incorporated models improve performance on existing benchmarks ImageNet-E and ImageNet-X and significantly enhance results on the ImageNet-PD dataset while maintaining consistent performance on standard data. Also, MPD incorporated in practical applications such as crowd counting, fisheye image recognition, and person re-identification improved performance. We believe that our work on mitigating perspective distortion could inspire future work to get new methods and address the challenges in other CV applications.
@@ 7 Perspectively distorted views synthesis through MPD
In connection to Fig. 1, we additionally show the non-black background variants of different orientations views synthesizing PD in Fig. 14.
Figure 14: MPD IB synthesizes perspective distortion with different orientations corresponding to the parameter. It replaces the black background with boundary foreground pixel values of the transformed image.
Fig. 15 Demonstrate controlled scaling of distortion in MPD-transformed image for different views in detail. It showcases the capabilities of MPD to mimic real-world distortions synthetically, which are otherwise challenging to capture to train the models.
@@ 9 MPD in self-supervised learning methods
We present a comprehensive investigation into the application of MPD (Geometry-Exploring Encoded Transformation) within self-supervised learning (SSL) approaches, with a primary focus on its integration with the contrastive learning method SimCLR [7], shown in Fig. 17. Our core objective is to assess MPD’s effectiveness in enhancing SSL’s representation learning capacities by adapting to perspective variations. Section 3.3 in the main paper explains the detailed results of MPD incorporation with simCLR contrastive learning to mitigate perspectivness. Further Table linear evaluation results are included in Section 12. Furthermore, we extend our exploration to evaluate MPD’s compatibility and performance in conjunction with another SSL method, DINO [6], which is grounded in knowledge distillation-based self-supervised approaches. Results on MPD incorporation in DINO (Fig. 18) are presented in Table 14 and Fig. 20 in Section 12.
Figure 17: ssl:MPD - MPD integrated in SimCLR [7] contrastive learning self-supervised method. MPD introduces perspectiveness by transforming the views of the input image. Standard SimCLR and ssl:DINO-MPD both are pretrained for 100 epochs with a batch size of 512 and linear evaluation for 100 epochs with a batch size of 256. ResNet50 is backbone.
Figure 18: ssl:DINO-MPD - MPD integrated in DINO [6] knowledge distillation based self-supervised method. MPD introduces perspectiveness by transforming the input image before the creation of views. Standard DINO and ssl:DINO-MPD both are pretrained for 100 epochs with a batch size of 512 and linear evaluation for 100 epochs with a batch size of 256. ViT-small is backbone.
@@
Extended evaluations on the ImageNet-PD benchmark dataset, detailed in Tables 6 and 7, reveal how different architectures respond to perspective distortions, extending the main paper’s insights (Fig. 4). Models like ResNet50 and VIT-L-16 show varied robustness to distortions such as Left-view (PD-L and PD-LI), with ResNet50 maintaining 63.37% Top-1 accuracy for PD-L compared to 61.15% for PD-B, and VIT-L-16 at 74.34% for PD-LI versus 71.73% for PD-BI. With their self-attention mechanisms, Vision Transformers demonstrate a smaller performance drop than traditional CNNs under perspective distortions, indicating their potential to maintain representational robustness. As discussed in our main paper, this analysis underscores the need for adaptive, perspective-aware models like MPD in real-world applications.
Table 6: Detailed results on the performance of standard CNN architectures (trained on ImageNet dataset) and transformer; evaluated on original validation set and benchmark ImageNet-PD subsets (without background interpolation). Results in continuation of Fig. 4 (first plot).
@@
Table 7: Detailed results on the performance of standard CNN architectures (trained on ImageNet dataset) and transformer; evaluated on original validation set and ImageNet-PD subsets (with background interpolation). Results in continuation of Fig. 4 (second plot).
@@
In the extended analysis presented in Tables 8 and 9, we observe a shift in the performance of MPD and MPD IB when applied at a probability of 1.0, approximately following the trends reported in the main paper up to a probability of 0.8 (Fig. 9). For MPD at P=1.0, a marginal decrease in Top-1 accuracy on the original ImageNet validation set (74.10% $\pm$ 0.03) compared to P=0.8 (76.34% $\pm$ 0.02), the robustness against perspective distortions on ImageNet-PD subsets is maintained, with accuracies between 71.57% $\pm$ 0.04 to 72.05% $\pm$ 0.04. This highlights a nuanced balance between handling distortions and achieving peak performance on standard datasets. Similarly, MPD IB at P=1.0 follows a comparable pattern. The Top-1 accuracy on the standard ImageNet set slightly reduces to 74.62% $\pm$ 0.03 from 75.63% $\pm$ 0.05 at P=0.8. However, its performance on the ImageNet-PD subsets, though marginally lower than MPD, remains consistent with Top-1 accuracies in the range of 68.23% $\pm$ 0.04 to 70.30% $\pm$ 0.03. These results emphasize the trade-offs involved in fully applying MPD and MPD IB and the importance of finding a balanced approach to optimize performance across both standard data and perspectiveness present in data.
ImageNet-E: In our extended analysis of MPD and MPD IB within a supervised approach (Table 10 and Fig. 6), we observe notable performance improvements in size change scenarios of ImageNet-E, indicative of enhanced perspectiveness robustness. These models, specifically in size-related subsets (full, 0.10, 0.08, 0.05), show a reduced drop in Top-1 accuracy and higher absolute accuracy than standard ResNet50. However, in background change subsets, MPD and MPD IB’s performance occasionally falls behind ResNet50, reflecting their focus on perspective distortions over background variations. This further aligns with the findings on MPD’s capabilities in perspective-centric scenarios.
@@
Table 8: Extended comparative analysis of MPD trained on supervised (supervised:MPD model) and self-supervised (ssl:MPD model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background, without background interpolation). The probability $P$ of applying MPD in model fine-tuning is in the first column. Added results for probability values 0.0 and 1.0 in continuation to Table 1.
@@
Table 9: Extended comparative analysis of MPD IB trained on supervised (supervised:MPD IB model) and self-supervised (ssl:MPD IB model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background, without background interpolation). The probability $P$ of applying MPD BI in model fine-tuning is in the first column.
@@
Figure 19: Detailed performance comparison on ImageNet-E [22]: Detailed results on Drop of Top1 accuracy under background changes, size changes, random position (rp), random direction (rd), and average over 11 subsets. Lower is better. Average reports the mean of all subsets. MPD variants outperforms standard ResNet50 on subsets which are affected with perspectiveness, e.g., size, position, and direction. Corresponding details on absolute accuracy measure are in Table 10. Detailed analysis in continuation of Fig. 6; Fig. 6 is given in main paper.
@@
SimCLR [7] self-supervised MPD models demonstrate notable effectiveness on ImageNet-E and ImageNet-X, detailed in Tables 10 and 11. On ImageNet-E, self-supervised MPD variants show a marked reduction in error ratios in size-related subsets, with error ratios like 1.44 for ’larger’ and 1.46 for ’smaller’, indicating a strong grasp of perspective challenges. In ImageNet-X, these models excel in handling ’object blocking’ (error ratio 1.61) and ’person blocking’ (error ratio 1.97) factors, achieving higher absolute accuracy compared to standard ResNet50, as illustrated in Fig. 8. The detailed analysis in Fig. 19 further emphasizes the models’ enhanced adaptability in self-supervised supervised learning, underlining their robustness in understanding complex visual scenarios.

### v2

@@ 3.2.1 Existing Benchmarks:
Both supervised:MPD and supervised:MPD IB models are trained on ImageNet dataset and evaluated on existing benchmarks and ImageNet-PD. ImageNet-X [23] and ImageNet-E [28]) considered as they are partially affected by perspective distortion. ImageNet-E[28] is a recent benchmark dataset specifically designed to evaluate the robustness of image classifiers concerning object attributes, including background settings, object sizes, spatial positioning, and orientation. ImageNet-X[23] introduces a set of sixteen human annotations focusing on attributes like pose, background, and lighting derived from the ImageNet.
Figure 6: ImageNet-E[28]: Drop of Top1 accuracy under background changes, size changes, random position (rp), random direction (rd), and average over 11 subsets. Lower is better. The mean is reported for size and background-related subsets. Average reports the mean of all subsets.
Figure 7: Activation maps (ImageNet-E). ’redbone’ from rd and ’peke’ from rp subset.
@@ 3.2.2 Benchmarking on ImageNet-PD:
Both supervised:MPD and supervised:MPD IB demonstrate notable improvements (Fig. 9 (a)-(b)) on ImageNet-PD subsets, achieving average Top-1 accuracy enhancements of 9.42% and 9.54% respectively, over the standard ResNet50 model.
Table 1: Comparative analysis of MPD trained on supervised (supervised:MPD model) and self-supervised (ssl:MPD model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background). The probability $P$ of applying MPD in model fine-tuning is in the first column. More results on probabilities 0.0 and 1.0 in Table 9 in suppl. material.
@@
Please refer to Table 1 and Table 10 for detailed comparison; Table 10 is in suppl. material.
@@
Concurrently, these models maintain performance on the original ImageNet validation set, highlighting their efficacy and adaptability amidst perspective distortion. The activation maps shown in Fig. 5 (row 2) further corroborate this outperforming trend. Ablation on the probability of applying MPD and MPD IB as a data augmentation (Fig. 9 (c)-(d)) exhibits consistent performance. MPD also compared with other popular augmentation methods (Mixup [58], Cutout [12], AugMix [17], and Pixmix [18]) on ImageNet-PD and ImageNet-E, consistently outperforming them. Detailed results are in the suppl. material (refer to Table 17).
@@ 3.3 MPD’s effects on self-supervised learning
MPD as data augmentation integrated in SimCLR [7] and DINO [6] (small ViT [13] backbone), as in Fig. 19 and 20 in suppl. material. Models are pre-trained and finetuned / liner-probed on ImageNet and evaluated on existing benchmarks and ImageNet-PD.
@@ 3.3.1 Existing Benchmarks:
ssl:MPD and ssl:MPD IB models demonstrate efficacy in handling perspective-related variations in ImageNet-E as shown in Fig. 6. Specifically, it shows a reduction in the accuracy drop for position (1.21%), direction (1.41%), and size (0.48%) changes compared to the standard ResNet50, indicating its robustness. On the ImageNet-X benchmark, ssl:MPD and ssl:MPD IB models outperform with an average accuracy of 62.99%, compared to the standard ResNet50’s 60.64%, and also show a reduced error ratio of 1.41 against the ResNet50’s 1.55 (refer Fig. 8). Notably, performance improvement is evident in managing variations in size and shape, showcasing its robustness against PD. Detailed results are in Table 12 in suppl. material.
@@ 3.3.2 Benchmarking on ImageNet-PD:
ssl:MPD model (Fig. 10 (a)) shows an average improvement of approximately 10.32% across the PD subsets, while the ssl:MPD IB (Fig. 10 (b)) demonstrates an average increase of around 10.14%. In addition, both variants display robust performance on the original ImageNet validation set. Please refer to Table 1 and Table 10 for detailed comparison in suppl. material. The activation maps depicted in Fig. 5 (row 3) further validate the improved performance. Moreover, ablation on the probability of applying MPD and MPD IB as data augmentation (Fig. 10 (c)-(d)) demonstrates a consistent increase in performance. We have also performed linear evaluations on self-supervised method SimCLR [7] and DINO [6] (refer Table 15, 16; Fig. 22 for detailed results in suppl. material). Similarly, significant improvement is noticed on ImageNet-PD subsets while maintaining the performance on standard ImageNet as shown in Fig. 11 (refer Table 13, 14 for detailed results in suppl. material).
@@ 3.4 MPD’s generalizability
This section discusses MPD’s adaptability in diverse applications, namely crowd counting (CC), fisheye, and person re-identification, and challenging CV tasks such as object detection, where perspective distortion is evident.
@@ 3.4.1 Crowd counting:
Numerous human heads within a single image present challenges due to the variations in size, orientation, and shape, naturally including perspective distortion. We investigate the potential of MPD in crowd counting on several publicly available datasets, including ShanghaiTech (SHTech) Part A and Part B[60] and UCF-CC50[21]. The ShanghaiTech Part A features highly congested scenes, and Part B consists of sparser scenes captured on a busy street. The UCF-CC50 dataset comprises 50 diverse scenes featuring a considerable variation in the crowd.
MPD-CC: MPD is adapted to CC by extending its capability to transform scene images and their CC labels, as mentioned in steps 4-7 of algorithm 2.
MPD-AutoCrowd: We introduce MPD-AutoCrowd (Algorithm 2), a novel approach for the automated augmentation of the crowd in scene images by upgrading MPD-CC. It seamlessly integrates synthetic crowds into the background regions of transformed images, complete with corresponding labels. This approach facilitates scene-level crowd augmentation without human supervision (Fig. 12).
@@
MPD-CC and MPD-AutoCrowd are integrated with P2P-Net[45] to evaluate the effectiveness of MPD in crowd counting. Training parameters, CNN architecture (VGG16), and evaluation configuration are used as mentioned in P2P-Net[45]. The probability of applying MPD-CC and MPD AutoCrowd is 0.1, and real and imaginary components of $c$ range from $0.2$ to $0.6$. This method has yielded improved results, surpassing previous state-of-the-art performance. Refer to Table 2 for a detailed comparison with other CC methods. MPD-CC and MPD-AutoCrowd outperform on multiple benchmarks.
MPD-AutoCrowd has a MAE of 50.81 on SHTech Part A and 6.73 on SHTech Part B. The performance on the UCF-CC50 ( a highly dense crowds) dataset, where MPD-AutoCrowd has an MAE of 96.80 and a Mean Squared Error (MSE) of 139.50, significantly outperforms other models. We further improved results on UCF-CC50 with ResNet50 encoder as shown in Table 18 in suppl. material. We also adapted other perspective transforms[59][38] and RandAug[10] augmentation for crowd counting and compared with MPD-CC and MPD-AutoCrowd which shows that our proposed methods outperform. Details are in suppl. material (refer Table 19 and Fig. 23 in sec. 14.1).
@@ 3.4.2 Transfer learning on fisheye images
Fisheye image distortions, caused by a wide field of view, are handled by spherical perspective models in [61, 27]. MPD simulates spherical views in pixel space and improves representation learning for fisheye images. We showcase the transferable capability of models trained with MPD and MPD IB through supervised and self-supervised methods on the ImageNet dataset, applying them to the VOC-360 fisheye dataset[14]. This dataset has 39,575 fisheye images with multi-label classification labels. We fine-tune these models on the VOC360, running for 100 epochs with a batch size 16. We show the performance of supervised and self-supervised models for the multi-label classification task in Table 3. ResNet50 is used as backbone. Ablations on label efficiency (Table 20, 21, and Fig. 24) are in suppl. material.
@@ 3.4.3 Person Re-Identification:
The DeepSportRadar re-identification dataset[48] includes video frames with varying poses and camera angles, offering training (436 queries, 8133 gallery images) and testing sets (50 queries, 910 gallery images) to address perspective variability in re-identification tasks. We incorporated the MPD within the state-of-the-art CLIP-ReIdent method[15]. The probability of applying MPD is 0.1, and real & imaginary components of $c$ range from $0.2$ to $0.6$. CLIP-ReIdent is CLIP[41]-adapted contrastive image-image pre-training for person re-identification, leveraging transformer architecture. MPD-incorporated Clip-ReIdent models improve performance with both backbone architectures, ViT-L-14 and ResNet50x16 (ref. Table 5). Ablations on MPD parameters are in suppl. material in Table 22. In addition, it boosts performance across different frame counts, achieving competitive results with limited frames and performance further enhanced by longer training (refer Table 5).
@@
3.4.4 Object Detection:
We applied MPD method for the object detection (MPD-OD) by transforming image and bounding boxes as augmentation to mimic PD (fig. 14). We have trained FasterRCNN [43] and FCOS [47] with MPD-OD on COCO [30]. Applying MPD-OD to 50% probability (P=0.5), we achieved 3% and 1.6% gain over original models (Table 6). Ablations are in suppl. material (Table 23 and Fig. 25 in sec. 14.4).
Method Original MPD-OD (IoU=0.50:0.95/IoU=0.50)
FasterRCNN[43] 37.10/55.80 40.00/61.10
FCOS[47] 38.60/57.40 40.20/60.30
Table 6: MPD-OD on COCO dataset.
Figure 14: MPD-OD examples
@@ 4 Conclusion
We address perspective distortion for robust computer vision applications. The proposed MPD method models real-world distortions by applying a family of Möbius transforms on existing data without estimating camera parameters or relying on actual distorted data. We develop a perspectively-distorted ImageNet-PD dataset to benchmark the robustness of the computer vision models and reveal that existing models lack robustness. Further, MPD-incorporated models improve performance on existing benchmarks ImageNet-E and ImageNet-X and significantly enhance results on the ImageNet-PD while maintaining consistent performance on standard data. Also, MPD incorporated in practical applications such as crowd counting, fisheye image recognition, person re-identification, and object detection improved performance. We believe that our work on mitigating PD could inspire future work for other CV challenges. Acknowledgment The authors thank Sumit Rakesh, Luleå University of Technology, for his support with the Lotty Bruzelius cluster. We also thank the National Supercomputer Centre at Linköping University for the Berzelius supercomputing, supported by the Knut and Alice Wallenberg Foundation.
@@ 7 Perspectively distorted views synthesis through MPD
MPD distorts the original image to synthesize perspective distortion based on steps shown in Algorithm 1 in main paper, resulting in a distorted image with a black background (variant 1, Fig. 15). Optionally, to remove the black background, replication padding is used to obtain integrated padding variant (variant 2, Fig. 15) of transformed image in step 9 of Algorithm 1.
Figure 15: MPD transformed images - (black background) variant 1 and (integrated padding background) variant 2
In connection to Fig. 1, we additionally show the integrated padding background variants of different orientations views synthesizing PD in Fig. 16.
Figure 16: MPD IB (integrated padding background) synthesizes perspective distortion with different orientations corresponding to the parameters. It replaces the black background with boundary foreground pixel values of the transformed image.
Fig. 17 Demonstrate controlled scaling of distortion in MPD-transformed image for different views in detail. It showcases the capabilities of MPD to mimic real-world distortions synthetically, which are otherwise challenging to capture and train the models.
@@ 9 MPD in self-supervised learning methods
We present a comprehensive investigation into the application of MPD (Geometry-Exploring Encoded Transformation) within self-supervised learning (SSL) approaches, with a primary focus on its integration with the contrastive learning method SimCLR [7], shown in Fig. 19. Our core objective is to assess MPD’s effectiveness in enhancing SSL’s representation learning capacities by adapting to perspective variations. Section 3.3 in the main paper explains the detailed results of MPD incorporation with simCLR contrastive learning to mitigate perspective distortion. Further Table linear evaluation results are included in Section 12. Furthermore, we extend our exploration to evaluate MPD’s compatibility and performance in conjunction with another SSL method, DINO [6], which is grounded in knowledge distillation-based self-supervised approaches. Results on MPD incorporation in DINO (Fig. 20) are presented in Table 15 and Fig. 22 in Section 12.
Figure 19: ssl:MPD - MPD integrated in SimCLR [7] contrastive learning self-supervised method. MPD introduces perspective distortion by transforming the views of the input image. Standard SimCLR and ssl:DINO-MPD both are pretrained for 100 epochs with a batch size of 512 and linear evaluation for 100 epochs with a batch size of 256. ResNet50 is backbone.
Figure 20: ssl:DINO-MPD - MPD integrated in DINO [6] knowledge distillation based self-supervised method. MPD introduces perspective distortion by transforming the input image before the creation of views. Standard DINO and ssl:DINO-MPD both are pretrained for 100 epochs with a batch size of 512 and linear evaluation for 100 epochs with a batch size of 256. ViT-small is backbone.
@@
Extended evaluations on the ImageNet-PD benchmark dataset, detailed in Tables 7 and 8, reveal how different architectures respond to perspective distortions, extending the main paper’s insights (Fig. 4). Models like ResNet50 and VIT-L-16 show varied robustness to distortions such as Left-view (PD-L and PD-LI), with ResNet50 maintaining 63.37% Top-1 accuracy for PD-L compared to 61.15% for PD-B, and VIT-L-16 at 74.34% for PD-LI versus 71.73% for PD-BI. With their self-attention mechanisms, Vision Transformers demonstrate a smaller performance drop than traditional CNNs under perspective distortions, indicating their potential to maintain representational robustness. As discussed in our main paper, this analysis underscores the need for adaptive, perspective-aware models like MPD in real-world applications.
Table 7: Detailed results of ImageNet trained models with standard CNN and transformer architectures; evaluated on original validation set and ImageNet-PD subsets with black background. Results in continuation of Fig. 4 (first plot).
@@
Table 8: Detailed results of ImageNet trained models with standard CNN and transformer architectures; evaluated on original validation set and ImageNet-PD subsets with integrated padding background. Results in continuation of Fig. 4 (second plot).
@@
In the extended analysis presented in Tables 9 and 10, we observe a shift in the performance of MPD and MPD IB when applied at a probability of 1.0, approximately following the trends reported in the main paper up to a probability of 0.8 (Fig. 9). For MPD at P=1.0, a marginal decrease in Top-1 accuracy on the original ImageNet validation set (74.10% $\pm$ 0.03) compared to P=0.8 (76.34% $\pm$ 0.02), the robustness against perspective distortions on ImageNet-PD subsets is maintained, with accuracies between 71.57% $\pm$ 0.04 to 72.05% $\pm$ 0.04. This highlights a nuanced balance between handling distortions and achieving peak performance on standard datasets. Similarly, MPD IB at P=1.0 follows a comparable pattern. The Top-1 accuracy on the standard ImageNet set slightly reduces to 74.62% $\pm$ 0.03 from 75.63% $\pm$ 0.05 at P=0.8. However, its performance on the ImageNet-PD subsets, though marginally lower than MPD, remains consistent with Top-1 accuracies in the range of 68.23% $\pm$ 0.04 to 70.30% $\pm$ 0.03. These results emphasize the trade-offs involved in fully applying MPD and MPD IB and the importance of finding a balanced approach to optimize performance across both standard data and perspective distortion present in data.
ImageNet-E: In our extended analysis of MPD and MPD IB within a supervised approach (Table 11 and Fig. 6), we observe notable performance improvements in size change scenarios of ImageNet-E, indicative of enhanced robustness against perspective distortion. These models, specifically in size-related subsets (full, 0.10, 0.08, 0.05), show a reduced drop in Top-1 accuracy and higher absolute accuracy than standard ResNet50. However, in background change subsets, MPD and MPD IB’s performance occasionally falls behind ResNet50, reflecting their focus on perspective distortions over background variations. This further aligns with the findings on MPD’s capabilities in perspective-centric scenarios.
@@
Table 9: Extended comparative analysis of MPD trained on supervised (supervised:MPD model) and self-supervised (ssl:MPD model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background). The probability $P$ of applying MPD in model fine-tuning is in the first column. Added results for probability values 0.0 and 1.0 in continuation to Table 1.
@@
Table 10: Extended comparative analysis of MPD IB trained on supervised (supervised:MPD IB model) and self-supervised (ssl:MPD IB model) approaches and evaluated on original ImageNet and ImageNet-PD subsets (black background). The probability $P$ of applying MPD BI in model fine-tuning is in the first column.
@@
Figure 21: Detailed performance comparison on ImageNet-E [28]: Detailed results on Drop of Top1 accuracy under background changes, size changes, random position (rp), random direction (rd), and average over 11 subsets. Lower is better. Average reports the mean of all subsets. MPD variants outperforms standard ResNet50 on subsets which are affected with perspective distortion, e.g., size, position, and direction. Corresponding details on absolute accuracy measure are in Table 11. Detailed analysis in continuation of Fig. 6; Fig. 6 is given in main paper.
@@
SimCLR [7] self-supervised MPD models demonstrate notable effectiveness on ImageNet-E and ImageNet-X, detailed in Tables 11 and 12. On ImageNet-E, self-supervised MPD variants show a marked reduction in error ratios in size-related subsets, with error ratios like 1.44 for ’larger’ and 1.46 for ’smaller’, indicating a strong grasp of perspective challenges. In ImageNet-X, these models excel in handling ’object blocking’ (error ratio 1.61) and ’person blocking’ (error ratio 1.97) factors, achieving higher absolute accuracy compared to standard ResNet50, as illustrated in Fig. 8. The detailed analysis in Fig. 21 further emphasizes the models’ enhanced adaptability in self-supervised supervised learning, underlining their robustness in understanding complex visual scenarios.